# Real-Time Wayfinding Assistant for Blind and Low-Vision Users

### v1

@@ Abstract.
Navigating unfamiliar places continues to be one of the most persistent and essential everyday obstacles for those who are blind or have limited vision (BLV). Existing assistive technologies, such as GPS-based navigation systems, AI-powered smart glasses, and sonar-equipped canes, often face limitations in real-time obstacle avoidance, precise localization, and adaptability to dynamic surroundings. To investigate potential solutions, we introduced PathFinder, a novel map-less navigation system that explores different models for understanding 2D images, including Vision Language Models (VLMs), Large Language Models (LLMs), and employs monocular depth estimation for free-path detection. Our approach integrates a Depth-First Search (DFS) algorithm on depth images to determine the longest obstacle-free path, ensuring optimal route selection while maintaining computational efficiency. We conducted comparative evaluations against existing AI-powered navigation methods and performed a usability study with BLV participants. The results demonstrate that PathFinder achieves a favorable balance between accuracy, computational efficiency, and real-time responsiveness. Notably, it reduces mean absolute error (MAE) and improves decision-making speed in outdoor navigation compared to AI-based alternatives. Participant feedback emphasizes the system’s usability and effectiveness in outside situations, but also identifies issues in complicated indoor locations and low-light conditions. Usability testing revealed that 73% of participants understood how to use the app in about a minute, and 80% praised its balance of accuracy, quick response, and overall convenience. By addressing critical usability and performance criteria through a low-cost, efficient, and scalable design, our study helps to advance AI-driven assistive technology for people with BLV, supporting greater independence and accessibility in daily navigation.
Blind and low vision (BLV) people, Indoor and outdoor navigation, Assistive technology, Large Language Models (LLMs), Depth estimation, Depth First Search (DFS), Computer Vision
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Human-centered computing Accessibility technologiesccs: Computing methodologies Computer vision problemsccs: Information systems Navigation-based services [^1] Code and dataset are available at: https://github.com/dabbrata/AI_Assisted_Nav_DFS and https://www.kaggle.com/datasets/artemmmtry/diode-a-dense-indoor-and-outdoor-depth-dataset
@@ 1. Introduction
Navigating unfamiliar environments is a significant challenge for blind and low-vision (BLV) individuals, both indoors and outdoors. While traditional mobility aids such as white canes and guide dogs provide essential support, they have limitations in detecting obstacles and identifying clear paths in complex environments. To enhance mobility, BLV individuals rely on assistive technologies such as GPS-based navigation systems, AI-powered smart glasses (Waisberg et al., 2024), and sonar-equipped smart canes (Messaoudi et al., 2020). These solutions provide audio or haptic feedback for obstacle detection and route guidance. However, despite these advancements, existing technologies still face limitations in real-time navigation, precise obstacle avoidance, and environmental awareness, particularly in dynamic and unfamiliar scenarios.
Many GPS-based assistive technologies, including Be My Eyes (Redaktion, 2023), Seeing AI (Microsoft, 2017), and Aira (Aira Tech Corp, 2015), offer voice guidance and real-time assistance. However, they often struggle with providing precise localization, real-time obstacle avoidance, and effective pathfinding in unstructured environments. Additionally, AI-powered solutions, such as computer vision-based smart glasses (AlSaid et al., 2019), require high computational power, making them costly and less accessible for many users. Given these challenges, there is a need for a more efficient, low-cost, and user-friendly navigation solution that leverages image processing and depth estimation techniques to assist BLV individuals effectively.
In this paper, we examine various navigation techniques, including those that use Vision Language Models (VLMs), and then present a computer vision-based algorithm for recognizing free space and enabling pathfinding. The system also incorporates monocular depth estimation to generate depth maps, which, when processed using LLMs, improve obstacle detection and free path identification. Finally, we propose a novel image-processing approach that leverages Depth-First Search (DFS) on depth images to identify the longest free path, ensuring optimal route selection for BLV individuals. This computationally efficient technique enhances real-time navigation by providing accurate obstacle-free paths while maintaining low processing costs. Compared to existing AI-powered solutions, our method offers a user-friendly, computationally efficient, and cost-effective alternative without compromising accuracy or responsiveness, making it a practical and accessible solution for BLV individuals to navigate safely and independently.
@@
  • RQ2: How effective are AI-powered solutions for real-time navigation compared to our method?
  • RQ3: How well does our method address the shortcomings of AI-powered solutions for real-time navigation?
  • RQ4: How well does the implemented system address BLV individuals’ needs, and what are their impressions of its effectiveness?
To answer RQ1, we conducted preliminary studies on BLV individuals’ navigation habits and identified key challenges, such as difficulties in anticipating environmental layouts, avoiding obstacles, and maintaining a straight path in dynamic settings. Our approach aims to overcome these challenges through a cost-effective, real-time navigation aid that can be used with a single-camera smartphone.
To answer RQ2, we compare our method with existing AI-powered solutions that rely on high-end computational models for navigation. While AI-based systems provide valuable insights, they often require significant resources and may not perform optimally in all environments. Our approach, utilizing depth maps and DFS-based free-path detection, offers a lightweight yet efficient alternative.
To answer RQ3, we analyze the primary limitations of current AI-powered navigation solutions, including high computational costs, reliance on expensive hardware, and potential inaccuracies in real-time obstacle detection. Our method addresses these issues by using a more accessible and computationally efficient approach, ensuring better accuracy, faster processing, and enhanced usability.
To answer RQ4, we are conducting user studies and surveys in February 2025 to evaluate how well the implemented system addresses BLV individuals’ needs and their impressions of its effectiveness. This evaluation aligns with ongoing advancements in assistive technology, urban infrastructure developments, and relevant policy changes. Early surveys assessing public interest in assistive navigation solutions have highlighted key concerns regarding privacy, data security, and long-term usability (Martin et al., 2010). Our study considers these factors to refine the system, enhance user trust, and improve adoption, ensuring it effectively meets the real-world navigation challenges faced by BLV individuals.
While our system offers significant advantages in terms of accessibility, efficiency, and accuracy, we acknowledge potential limitations related to privacy concerns and adaptability in highly dynamic environments. Our research adds to the larger goal of increasing accessibility and independence by creating a user-friendly, cost-effective, and efficient navigation system for BLV persons. Furthermore, the approaches investigated in this study can contribute to future advancements in real-time assistive technology for people with visual impairments, as well as inspire breakthroughs in autonomous navigation systems for personal and public use.
@@ 2. Related Works
Developing successful navigation systems for blind and low-vision (BLV) people has been a major focus of assistive technology research. Various methodologies have been investigated, ranging from GPS-based solutions and depth-sensing techniques to AI-powered systems and infrastructure-free navigation tools. While many of these technologies strive to increase mobility, obstacle detection, and spatial awareness, they frequently confront hurdles in real-time processing, environmental adaptability, and user accessibility. This section examines existing research on BLV navigation systems, depth estimate methods, and AI-powered assistive technologies, emphasizing their contributions, limits, and implications for the development of our proposed solution. Table 1 highlights significant studies on BLV navigation and accessibility, including techniques, conclusions, and limitations.
@@ 2.1. Navigation Systems for BLV Individuals
Navigating unknown regions poses significant obstacles for blind and low-vision (BLV) people, demanding intensive study into assisted navigation technologies. A variety of approaches have been investigated, ranging from GPS-based solutions to sensor-driven indoor navigation systems.
Traditional GPS-based technology have been important in assisting BLV patients. Microsoft Soundscape (Rickert, 2010), for example, improves spatial awareness with 3D audio cues. Liu et al. (Liu et al., 2022) studied user interaction with Soundscape and discovered that spatial audio callouts considerably improved usability and retention. Their investigation did, however, identify limits in exact localization and real-time obstacle avoidance, both of which are significant challenges in GPS-based systems.
To address the limitations of GPS-based technology, other solutions based on ambient sensors and computer vision have been investigated. Jain et al. (Jain et al., 2023) describes a system that uses current street cameras to deliver real-time aural feedback to BLV pedestrians. By using computer vision algorithms to street camera feeds, the system provides users with detailed spatial information that aids in obstacle avoidance and safe street crossing. While this strategy improves real-time navigation, its reliance on public monitoring infrastructure limits its scalability.
For indoor navigation, Tsai et al. (Tsai et al., 2024) proposed inertial-based iOS applications—Wayfinding and Backtracking—that use a smartphone’s internal sensors for pathfinding without relying on external infrastructure like Bluetooth beacons. The Wayfinding software enables hands-free navigation along established routes, while Backtracking allows users to retrace their steps in unfamiliar areas. Although these systems are useful, they suffer from localization drift—a progressive rise in positioning errors over time—which leads to lower navigation accuracy, especially in bigger or featureless areas.
Depth sensing combined into portable devices has also emerged as a viable solution. See et al. (See et al., 2022) created a smartphone application that uses depth cameras to identify obstacles and recognize objects. This infrastructure-free technique provides real-time environmental awareness via a mobile device. In controlled experiments, the system detected obstacles accurately, although performance varied depending on environmental conditions such as lighting and item size.
Robotic help expands the capabilities of navigation aids. Kuribayashi et al. (Kuribayashi et al., 2023) built a suitcase-shaped robot (PathFinder) that helps blind people navigate unfamiliar indoor places by dynamically recognizing junctions and reading signs. The solution gives real-time input without relying on prebuilt maps, giving users more flexibility. However, it needs users to recall routes while navigating, which might raise cognitive burden and impair usefulness.
Ren et al. (Ren et al., 2023) proposed RouteNav, which is a hybrid system that combines GPS localization and inertial dead reckoning to increase navigation accuracy in complex areas like transportation hubs. This system mitigates the impact of GPS inaccuracy by segmenting environments into a tile-based navigation system that provides multimodal feedback such as voice, haptics, and aural signals. While the system is excellent in organized surroundings, it still faces issues with localization drift and GPS signal discrepancies, particularly in underground or open locations.
Table 1. Summary of Related Works on BLV Navigation and Accessibility
Author(s) Focus Area Methodology Key Findings Limitations
Liu et al. [6] GPS-based outdoor navigation Uses spatial audio callouts for location awareness Improves spatial perception and usability Limited accuracy in localization and obstacle avoidance
Tsai et al. [11] Indoor navigation without external infrastructure Uses smartphone sensors for path tracking Enables retracing steps and predefined routes Suffers from localization drift over time
See et al. [8] Depth camera-based obstacle detection Uses embedded depth cameras for feedback Provides auditory/haptic feedback for real-time navigation Performance varies by environment, limited object categories (e.g., furniture, doorways)
Seo et al. [9] AI-driven data visualization for BLV users Combines text, audio, and haptic feedback Enhances structured interaction with complex data AI hallucinations, difficulty in prompt engineering
Zhang et al. [12] Accessible 2D data visualization for BLV users Multimodal interfaces (touch, audio, haptic) Supports scatter plots, bar charts, and line charts Complex gestures, small sample sizes in testing
Jain et al. [9] Real-time urban navigation using public cameras Analyzes live video feeds for path guidance Adapts to dynamic obstacles in cities Requires permanent infrastructure, limiting scalability
Shaheen [10] Scalability in educational accessibility Assesses assistive tech usage in education Highlights lack of BLV-friendly resources Requires external assistive devices for basic access
@@ 2.2. AI-Powered Scene Description and Accessibility Tools
Beyond navigation, AI-powered scene description tools have played an important role in increasing BLV users’ environmental awareness. These applications combine computer vision and natural language processing to deliver real-time descriptions of the environment, objects, and people in order to promote user independence. However, issues such as trust in AI outputs, description accuracy, and user experience remain major concerns.
Penuela et al. (Gonzalez Penuela et al., 2024) looked into how BLV users engage with AI-powered scene description apps, including their trust, satisfaction, and usage habits. The study discovered that users regularly used these tools for object recognition and scene assessment in familiar interior settings, based on a two-week diary study with 16 participants. Interestingly, despite occasional mistakes in AI-generated descriptions, users frequently relied on the system, demonstrating a nuanced link between trust and perceived utility. The study also stressed the significance of increasing AI accuracy and user customisation choices in order to improve satisfaction and usability.
Another interesting method (See et al., 2022) is to include depth photography into assistive applications to improve obstacle identification and spatial awareness (See et al., 2022). A depth camera-based navigation tool was created to help users recognize barriers and locate things by providing aural and haptic input, removing the need for extra external hardware. The technology successfully supplied real-time input but had limitations in different contexts. Furthermore, the object detection model was limited to specific categories, such as furniture, entrances, and everyday domestic things, restricting its applicability in unknown environments.
To broaden the scope of AI-powered accessibility, Seo et al. (Seo et al., 2024) explored how BLV users interact with multimodal large language models (LLMs) to interpret complicated data visualizations. The study launched maidrAI, an AI-powered data visualization tool that combines text, audio, and haptic input, allowing BLV users to interact with data using organized, customized descriptions. User research found that participants preferred clear, organized outputs and valued the opportunity to tailor AI-generated solutions. However, issues such as AI hallucinations (inaccurate or falsified information) and the difficulty of rapid engineering remained, emphasizing the need for more dependable and user-friendly AI systems.
Efforts to increase data accessibility for BLV users, ChartA11y(Zhang et al., 2024) allows blind people to interact with 2D data visualizations via multimodal interfaces on their smartphones. The technology, which combines touch-based exploration with audio and haptic feedback, gives users access to scatter plots, bar charts, and line charts. User tests demonstrated the usefulness of mixing structured navigation with direct touch interactions, however obstacles such as gesture complexity and small sample sizes in testing indicated areas for additional research.
@@ 2.3. User Engagement and Broader Assistive Technologies
User-centered design and engagement strategies are crucial for developing assistive technologies that meet the requirements and preferences of blind and low-vision (BLV) people. Several studies have looked into how design decisions affect user adoption, satisfaction, and long-term involvement with navigation and accessibility technologies.
In their study, Liu et al. (Liu et al., 2022) examined user behavior using Microsoft Soundscape to better understand the aspects that influence long-term engagement. The study found that early user interactions, particularly those during the first week, were major indicators of continuing use. Features such as spatial audio callouts and street previews helped to retain users. However, the study also noted difficulties in extrapolating findings to other navigation tools and advocated for the incorporation of qualitative user feedback to better customize future solutions.
Tsai et al. (Tsai et al., 2024) highlight the need for flexible, infrastructure-free navigation tools by introducing iOS applications for interior navigation using a phone’s built-in inertial sensors. By eliminating the requirement for external beacons or pre-installed infrastructure, the system enabled hands-free navigation, which is critical for BLV people who frequently rely on canes or guide dogs. While user studies validated the system’s usefulness, concerns like as localization drift and decreased accuracy in wide, open settings hampered its general applicability.
Educational accessibility for BLV students has also received substantial attention in assistive technology research. Shaheen (2024) (Shaheen, 2024) investigated the difficulties encountered by BLV students in K-12 school, particularly in using instructional tools. The study discovered that many children gained assistive technology (AT) literacy outside of school due to inadequacies in formal education assistance. Complex user interfaces, unlabeled buttons, and irregular navigation frameworks hampered academic involvement, particularly in STEM disciplines.
Urban navigation provides new issues, notably in picking the most relevant information to assist BLV users during mobility activities. Hoogsteen et al. (Hoogsteen et al., 2022) explored which environmental characteristics are easiest to describe for blind pedestrians. Intersections, traffic lights, and crosswalks were judged to be the most useful, while things such as poles or fire hydrants were deemed less relevant. Interestingly, the study found that early and late blind people had varied information needs, with late blind users requiring more extensive environmental descriptions. The study provided a structured framework for selecting information in assistive navigation systems, with the goal of lowering cognitive burden and improving spatial awareness.

### v2

# Real-Time Assistive Navigation for the Visually Impaired: A Scalable Approach for Indoor and Outdoor Mobility

@@ Abstract.
Navigating unfamiliar environments remains one of the most persistent and critical challenges for people who are blind or have limited vision (BLV). Existing assistive tools often rely on online services or APIs, making them costly, internet-dependent, and less reliable in real-time use. To address these limitations, we propose PathFinder, a novel mapless mobile phone-based navigation system that operates fully offline. Our method processes monocular depth images and applies an efficient pathfinding algorithm to identify the longest, clearest obstacle-free route, ensuring optimal navigation with low computational cost. Comparative evaluations show that PathFinder reduces mean absolute error (MAE), speeds decision-making, and achieves real-time responsiveness indoors and outdoors. A usability study with 15 BLV participants confirmed its practicality, where 73% learned to operate it in under a minute, and 80% praised its accuracy, responsiveness, and convenience. Despite challenges in complex indoor layouts and low light, PathFinder offers a low-cost, scalable, reliable alternative.
Blind and low vision (BLV) people, Indoor and outdoor navigation, Assistive technology, Large Language Models (LLMs), Pathfinder Algorithm, Computer Vision
$c$s: Human-centered computing Accessibility technologiesccs: Computing methodologies Computer vision problemsccs: Information systems Navigation-based services
@@ 1. Introduction
Navigating unfamiliar environments is a significant challenge for blind and low-vision (BLV) individuals, both indoors and outdoors. While traditional mobility aids such as white canes and guide dogs provide essential support, these have limitations in detecting obstacles and identifying clear paths in complex environments. To enhance mobility, BLV individuals rely on assistive technologies such as GPS-based navigation systems, AI-powered smart glasses (Waisberg et al., 2024), and sonar-equipped smart canes (Messaoudi et al., 2020). These solutions provide audio or haptic feedback for obstacle detection and route guidance. However, despite these advancements, existing technologies (Bouwer et al., 2011) still face limitations in real-time navigation, precise obstacle avoidance, and environmental awareness, particularly in dynamic and unfamiliar scenarios.
Many GPS-based assistive technologies, including Be My Eyes (Redaktion, 2023), Seeing AI (Microsoft, 2017), and Aira (Aira Tech Corp, 2015), offer voice guidance (Karunaratna et al., 2024) and real-time assistance. However, they often struggle with providing precise localization, real-time obstacle avoidance, and effective pathfinding in unstructured environments. Additionally, AI-powered solutions, such as computer vision-based smart glasses (AlSaid et al., 2019), require high computational power, making them costly and less accessible for many users. Given these challenges, there is a need for a more efficient, low-cost, and user-friendly navigation solution that leverages image processing and depth estimation techniques to assist BLV individuals effectively.
In this paper, we examine various navigation techniques, including those that use Vision Language Models (VLMs), and then present a computer vision-based algorithm to aid in BLV navigation. Our proposed system incorporates monocular depth estimation to generate depth maps, which, when processed using LLMs, improve obstacle detection and free path identification. Finally, we propose a novel image-processing approach that applies an innovative pathfinding algorithm on depth images to identify the longest free path, ensuring optimal route selection from mobile phone camera. This computationally efficient technique enhances real-time navigation by providing accurate obstacle-free paths while maintaining low processing costs.
@@
  • RQ2: How effective are AI-powered solutions for real-time navigation, and how does our method address their shortcomings?
  • RQ3: How well does the implemented system address BLV individuals’ needs, and what are their impressions of its effectiveness?
To answer RQ1, we conducted a qualitative analysis of our mobile application (Section 5 ) with BLV individuals to explore why they would rather choose our proposed navigation system over existing methods, such as white canes, guide dogs, or standard GPS-based applications. The findings revealed recurring challenges in current practices, including difficulties in anticipating environmental layouts, reliably detecting obstacles, and maintaining orientation in dynamic settings. Participants emphasized the need for a tool that not only enhances safety but also supports independence without requiring specialized hardware or extensive training. These insights guided the design of our system, which leverages a single-camera smartphone to provide an accessible, user-friendly, and real-time navigation solution.
To answer RQ2, we performed a quantitative comparison (Section 4) between our method and existing AI-powered navigation solutions. While AI-based systems can assist navigation, results show that they do not yield significant improvements in navigation performance relative to their computational and financial costs. Many of these solutions depend on cloud-based APIs, which incur recurring expenses, introduce latency, and limit scalability for widespread adoption. In contrast, our system is designed to operate fully offline, eliminating the need for internet access, thereby ensuring scalability across diverse contexts and making it more affordable and reliable for BLV users globally. By integrating depth maps with a novel pathfinding algorithm, our approach improves accuracy and reduces processing time, offering a more practical alternative to resource-intensive AI solutions.
To answer RQ3, we have conducted user studies and surveys (Section 5) to assess our system’s effectiveness, usability, and long-term adoption potential. Early results indicate that BLV individuals value the system’s independence from constant connectivity and its ability to address real-world navigation challenges, such as dynamic obstacles and unfamiliar environments. In line with broader developments in assistive technology and policy, we also account for concerns around privacy, data security, and adaptability. Incorporating these considerations strengthens user trust and ensures the system’s relevance for long-term use.
While our system demonstrates notable strengths in accessibility, efficiency, and scalability, we recognize potential limitations related to adaptability in highly dynamic environments and the need for continuous user feedback.
@@ 2. Related Works
Developing successful navigation systems for blind and low-vision (BLV) people has been a major focus of assistive technology research. Various methodologies have been investigated, ranging from GPS-based solutions and depth-sensing techniques to AI-powered systems and infrastructure-free navigation tools. While many of these technologies strive to increase mobility, obstacle detection, and spatial awareness, they frequently confront hurdles in real-time processing, environmental adaptability, and user accessibility. This section examines existing research on BLV navigation systems, depth estimate methods, and AI-powered assistive technologies, emphasizing their contributions, limits, and implications for the development of our proposed solution.
@@ 2.1. Navigation Systems for BLV Individuals
Initiatives into developing effective navigation aid for BLV individuals have evolved over time, from GPS-based solutions to infrastructure-free mobile applications and robotic aids, each targeting a unique facet of mobility and environmental awareness.
Early GPS-based systems, like Microsoft Soundscape (Rickert, 2010), intended to improve spatial awareness by providing 3D audio cues. Liu et al. (Liu et al., 2022) discovered that spatial audio callouts increased usability and retention, but also revealed difficulties in precise localization and real-time obstacle avoidance. These disadvantages have prompted academics to investigate alternative ways to standard GPS.
One approach has been to utilize existing infrastructure and sensing capabilities. Research conducted by Jain et al. (Jain et al., 2023a) showed that street cameras can give real-time audio feedback, helping BLV pedestrians avoid barriers and cross roadways safely. While promising, this strategy is primarily reliant on public monitoring infrastructure, posing questions about scalability and adoption. Wayfinding and Backtracking apps developed by Tsai et al.’s (Tsai et al., 2024), use smartphone inertial sensors to facilitate pathfinding and retracing paths, are examples of infrastructure-free solutions for indoor environments. These solutions eliminate the requirement for bluetooth beacons or any fixed infrastructure, but they are prone to localization drift, particularly in huge or featureless areas.
Other researchers have investigated depth sensing and embodied systems. See et al. (See et al., 2022) created a smartphone app that utilizes built-in depth cameras to identify obstacles and recognize objects, providing real-time environmental awareness without specialist hardware. Kuribayashi et al. (Kuribayashi et al., 2023) developed PathFinder, a suitcase-shaped robotic assistant that recognizes junctions and signage to direct users indoors. While these approaches boost accessibility, they also present trade-offs, such as performance unpredictability in diverse situations and greater cognitive burden due to route remembering. Similarly, Ren et al. (Ren et al., 2023) built RouteNav, a hybrid system integrating GPS localization with inertial dead reckoning, which increased accuracy in structured areas like transportation hubs but still struggled with drift and signal discrepancies.
CrossingGuard (Guy and Truong, 2012) has shown that providing more intersection-level information (e.g., traffic signals, road width, and intersection geometry) greatly boosted BLV pedestrians’ confidence when crossing streets. The study also proved the viability of scaling such information via crowdsourcing. More contemporary systems, such as StreetNav (Jain et al., 2024), reuse existing street cameras to give exact outdoor navigation. StreetNav reduced swerving, led participants closer to their destinations, and addressed privacy concerns by relying on existing infrastructure. Together, these works show a path from early intersection-focused systems to modern infrastructure-based precision, emphasizing both the benefits and remaining constraints of scale, environmental variability, and occlusion in outdoor navigation.
@@ 2.2. AI-Powered Scene Description and Accessibility Tools
AI-powered scene description systems have been a key factor in improving BLV users’ awareness of their surroundings. These computer vision and NLP-based solutions seek to promote independence by delivering real-time descriptions of objects, people, and situations. However, trust, accuracy, and usability remain critical consideration factors.
Penuela et al. (Gonzalez Penuela et al., 2024) examined how BLV users engage with such applications, showing that while participants relied on AI-generated descriptions in daily life, errors and inconsistencies created tension between trust and perceived utility. The study emphasized the importance of improving accuracy and supporting customization to foster long-term adoption. Likewise, See et al. (See et al., 2022) demonstrated how depth cameras can provide real-time obstacle awareness through auditory and haptic feedback without requiring additional hardware, though performance varied across contexts and object categories.
More recent research has expanded AI accessibility to areas such as data visualization. Seo et al. (Seo et al., 2024) developed maidrAI, a multimodal system that combines text, audio, and haptics to assist BLV users in understanding complex charts. Participants praised clear, structured outputs and personalization elements, while issues such as AI hallucinations and prompt complexity restricted usage. ChartA11y (Zhang et al., 2024) allows users to engage with scatter plots, bar charts, and line charts through multimodal smartphone interfaces in a related effort. User tests demonstrated the promise of touch-based exploration mixed with aural feedback, but gesture complexity and small-scale evaluations revealed areas for improvement.
This collection of works demonstrates both the promise and the problems of AI-powered accessibility aids. While these technologies can increase independence, these solutions face recurring challenges with trust, reliability, and user control that must be solved before widespread real-world adoption.
@@ 2.3. User Engagement and Broader Assistive Technologies
Designing effective navigation and accessibility technologies needs more than just technical accuracy; it also depends on how people embrace, interact with, and use these systems over time. A increasing body of research highlights the value of user-centered design, personalization, and consideration of the larger social and educational contexts in which blind and low-vision (BLV) individuals use assistive technology.
Liu et al. (Liu et al., 2022) investigated user behavior with Microsoft Soundscape to find parameters influencing prolonged engagement. Their findings revealed that early experiences, such as the first week of use, were highly predictive of long-term adoption. Features like spatial audio callouts and previews were crucial in keeping consumers interested. Tsai et al. (Tsai et al., 2024) highlighted the importance of flexible, infrastructure-free navigation tools. Their Wayfinding and Backtracking programs enabled hands-free mobility without the need for extra hardware; however, limitations such as drift limited their general usefulness.
Aside from turn-by-turn navigation, researchers have proposed that assistance technologies should promote exploration and spatial learning. According to Jain et al.’s study (Jain et al., 2023b), BLV users frequently seek higher-level spatial knowledge, such as the overall shape and structure of places, to construct cognitive maps, rather than depending exclusively on directions. The study also found that preferences for information varied across individuals based on blindness onset, sociability, and mobility aid use, emphasizing the importance of customisation. It emphasized the collaborative element of exploration, in which BLV users routinely interact with trainers, peers, and spectators. This viewpoint reframes navigation systems as components of a larger ecology of interconnectedness, rather than simply tools for independence.
Research has also demonstrated how expanded accessibility tools influence learning and inclusiveness. Shaheen (Shaheen, 2024) examined BLV students’ problems in K-12 schools and discovered that many obtained assistive technology (AT) literacy outside of school due to lack of formal education assistance. Complex interfaces and unlabeled buttons were particularly detrimental to STEM student participation. Hoogsteen et al. (Hoogsteen et al., 2022) found that intersections, traffic signals, and crosswalks were highly desired for urban navigation. Information needs differed between early-blind and late-blind users. These findings highlight that content selection, user training, and long-term adoption are just as important as technical innovation.
Taken together, this body of studies demonstrates that good assistive technology cannot be evaluated solely on technological parameters. Early user experiences, support for exploration and cognitive mapping, sensitivity to social collaboration, and integration into larger educational and urban contexts are all critical for long-term adoption.